OpenStudy (anonymous):

Solve by the substitution method: 7m+n=24 m-2n=12

OpenStudy (anonymous):

m-2n=12 m=2n+12 Plug that m value into other equation

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

You have two sets of equations 1)7m+n=24 2)m-2n=12

OpenStudy (anonymous):

Your goal is to isolate a variable one equation, so 2) m-2n=12 m=2n+12 This tells us that m is same as 2n+12 which means whenever we see 'm', we can instead write 2n+12. Plugging that into first equation 7m+n=24 We know what m is 7(2n+12)+n=24 14n+84=24 you can solve for n from here

OpenStudy (anonymous):

So M then equals 24? So on the other equation, I would say 24-2n=12?

OpenStudy (anonymous):

24-2n=12 so n= 6?

OpenStudy (anonymous):

where are you getting 24-2n=12 from?

OpenStudy (anonymous):

Well I thought that is m=24 I would plug 24 in to M for the second equation.

OpenStudy (anonymous):

how is m=24?